OpenStudy (anonymous):

Solve. 5x^2 - 3x = (5x^2 - 7x) + (8x - 2)

OpenStudy (anonymous):

A. x = -2 B. x = 2 C. x = -1/2 D. x = 1/2

OpenStudy (anonymous):

collect like terms on the right hand side

OpenStudy (anonymous):

remove the parenthesis first to remove the sign of grouping. .

OpenStudy (anonymous):

How am I supposed to set this up? :(

OpenStudy (mertsj):

\[5x ^{2}-3x=5x ^{2}-7x+8x-2\]

OpenStudy (anonymous):

after removing the parenthesis, group like terms. in grouping like terms, add if the same sign. subtract if of opposite sign. . .

OpenStudy (mertsj):

Subtract 5x^2 from both sides and get: \[-3x=x-2\]

OpenStudy (mertsj):

Subtract x from both sides and then divide by -4